OpenStudy (anonymous):

(10x^2)(6x^3) step by step please??

OpenStudy (saifoo.khan):

60x^5

OpenStudy (anonymous):

(10x^2)(6x^3)= (10)(x^2)(6)(x^3)= (10)(6)(x^2)(x^3)= (10*6)(x^(3+2)) 60x^5

OpenStudy (anonymous):

60x^(2+3) 60X^5

OpenStudy (anonymous):

thanks guys